# Optimal convergence rates in trace distance and relative entropy for the quantum central limit theorem

### v1

@@ 7 Proof of Theorem 2
Let $\sigma_{n}$, $\tilde{\sigma}_{n}$, and $\tilde{\sigma}_{n,G}$ be the same operators as defined in Section 4. The first point here is that, using Lemma 7 with $s=5$, we can write
$\big{\|}\rho^{\boxplus n}-\sigma_{n}^{\boxplus n}\big{\|}_{1}=\mathcal{O}\Big{(}\frac{1}{n^{(2+\delta)/2}}\Big{)},$
which in turn, using $\tilde{\sigma}_{n}^{\boxplus n}=D^{\dagger}_{z_{n}\sqrt{n}}\sigma_{n}^{\boxplus n}D_{z_{n}\sqrt{n}}$ and Lemma 8, implies
@@
Also, using Lemma 9 with $s=5$, we have
@@
where the sum is over all tuples $\mu,\nu$ of non-negative integers satisfying $\sum_{j}(\mu_{j}+\nu_{j})\leq 2(m+3)$, the $b_{\mu,\nu}$ are some constants. Thus, it suffices to verify that
@@
where $C=C(\nu_{\min},\nu_{\max})>0$ is the constant in Lemma 5 associated with the state $\rho$. Also, let $\epsilon\geq\epsilon_{n}$ be a constant to be determined. By using Plancherel identity, we can write
@@
where $Q(\cdot)$ is a polynomial with coefficients depending only on the covariance matrix of $\rho$. Now, using Lemma 10 with $s=5$, it is verified that there is a constant $C_{1}$ such that
@@
Putting all these bounds in (87), once again there exist polynomials $L_{1}(\cdot)$ and $L_{2}(\cdot)$, such that for sufficiently large $n$ and any $\|z\|\leq\epsilon_{n}\sqrt{n}$, we have

### v2

@@ 7 Proof of Theorem 2
Let $\sigma_{n}$, $\tilde{\sigma}_{n}$, and $\tilde{\sigma}_{n,G}$ be the same operators as defined in Section 4. Applying triangle’s inequality, we have
$\big{\|}\rho^{\boxplus n}-\tilde{\sigma}_{n}^{\boxplus n}\big{\|}_{1}$ $\leq\big{\|}\rho^{\boxplus n}-\sigma_{n}^{\boxplus n}\big{\|}_{1}+\big{\|}\sigma_{n}^{\boxplus n}-\sigma_{n,G}\big{\|}_{1}+\big{\|}\sigma_{n,G}-\tilde{\sigma}_{n,G}\big{\|}_{1}+\big{\|}\tilde{\sigma}_{n}^{\boxplus n}-\tilde{\sigma}_{n,G}\big{\|}_{1}$
$=\big{\|}\rho^{\boxplus n}-\sigma_{n}^{\boxplus n}\big{\|}_{1}+\big{\|}\sigma_{n,G}-\tilde{\sigma}_{n,G}\big{\|}_{1}+2\big{\|}\tilde{\sigma}_{n}^{\boxplus n}-\tilde{\sigma}_{n,G}\big{\|}_{1}.$
Then, using Lemma 7 with $s=4+\delta$, we find that the first term is bounded by $\mathcal{O}\big{(}n^{-(2+\delta)/2}\big{)}$. Also, applying similar computations as in Claim II and Claim III in Section 6, but with the stronger condition of finiteness of moments of order $4+\delta$, we find that the other two terms are also bounded by $\mathcal{O}\big{(}n^{-(2+\delta)/2}\big{)}$. Therefore,
@@
Moreover, using Lemma 9 with $s=4+\delta$, we have
@@
where the sum is over all tuples $\mu,\nu$ of non-negative integers satisfying $\sum_{j}(\mu_{j}+\nu_{j})\leq 2(m+3)$, and $b_{\mu,\nu}$’s are some constants. Thus, it suffices to verify that
@@
where $C=C(\nu_{\min},\nu_{\max})>0$ is the constant in Lemma 5 associated with the state $\rho$. Also, let $\epsilon\geq\epsilon_{n}$ be a constant to be determined. By using Plancherel’s identity, we can write
@@
where $Q(\cdot)$ is a polynomial with coefficients depending only on the covariance matrix of $\rho$. Now, using Lemma 10 with $s=4+\delta$, it is verified that there is a constant $C_{1}$ such that
@@
Putting all these bounds in (88), once again there exist polynomials $L_{1}(\cdot)$ and $L_{2}(\cdot)$ such that for sufficiently large $n$ and any $\|z\|\leq\epsilon_{n}\sqrt{n}$, we have